OpenStudy (anonymous):

Find the fifth roots of 32(cos 280° + i sin 280° I know i need to use DeMoivre's theorem but i need a refresher on how to do the process, I know that at some point all you do is add a certain amount into the equation too.

OpenStudy (anonymous):

What is the DeMoivre's theorem? Can you state it out?

OpenStudy (anonymous):

@e.mccormick @hartnn

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8

OpenStudy (amistre64):

http://www.mathwords.com/d/demoivres_thm.htm

OpenStudy (anonymous):

Thank you do you know if I would make the 32 into 2^5 ?

OpenStudy (amistre64):

that might be useful yes [32(cos 280° + i sin 280°)]^(1/5) 2(cos 280/5° + i sin 280/5°)

OpenStudy (anonymous):

And don't I add 360 to the angles? or do i also have to divide 360 by 5 for 72?

OpenStudy (amistre64):

you can do \[\frac{\theta+360k}{n}\]yes

OpenStudy (anonymous):

where k are the roots, 0, 1, 2, 3, 4? What would n be?

OpenStudy (anonymous):

5?

OpenStudy (amistre64):

n = 5 ...

OpenStudy (anonymous):

okay, sorry,

OpenStudy (amistre64):

a fifth root is when the exponent is 1/5 so the angle is multiplied by the 1/5 ... or simply a divide by 5 in this case

OpenStudy (anonymous):

The theta would be 280, or after dividing my 5, 56, correct?

OpenStudy (amistre64):

360/5 = 72 so the angles are going to start at 280 and increase/decrease by 72 is anothe rway to look at it

OpenStudy (anonymous):

okay, thank you very much @amistre64

OpenStudy (amistre64):

your welcome